### v1

# StreamBridge: Turning Your Offline Video Large Language Model into a Proactive Streaming Assistant

Abstract
We present StreamBridge, a simple yet effective framework that seamlessly transforms offline Video-LLMs into streaming-capable models. It addresses two fundamental challenges in adapting existing models into online scenarios: (1) limited capability for multi-turn real-time understanding, and (2) lack of proactive response mechanisms. Specifically, StreamBridge incorporates (1) a memory buffer combined with a round-decayed compression strategy, supporting long-context multi-turn interactions, and (2) a decoupled, lightweight activation model that can be effortlessly integrated into existing Video-LLMs, enabling continuous proactive responses. To further support StreamBridge, we construct Stream-IT, a large-scale dataset tailored for streaming video understanding, featuring interleaved video-text sequences and diverse instruction formats. Extensive experiments show that StreamBridge significantly improves the streaming understanding capabilities of offline Video-LLMs across various tasks, outperforming even proprietary models such as GPT-4o and Gemini 1.5 Pro. Simultaneously, it achieves competitive or superior performance on standard video understanding benchmarks.
1 Introduction
Video Large Language Models (Video-LLMs) [1; 2; 3; 4; 5] typically process entire pre-recorded videos at once. However, emerging applications, such as robotics [6; 7] and autonomous driving [8; 9], require causal perception and interpretation of visual information online. This fundamental mismatch highlights a critical limitation of current Video-LLMs, as they are not inherently equipped to operate in streaming scenarios where timely understanding and responsiveness are paramount.
Figure 1 highlights two representative patterns in streaming video understanding, which also correspond to the key challenges in adapting Video-LLMs from offline to streaming scenarios: (1) multi-turn real-time understanding and (2) proactive response generation. The first pattern involves multi-turn interactions, where the assistant receives user queries at different timestamps. In each turn, while keeping accumulated visual and conversational context as historical information, the model should focus on the most recent video segment. The second pattern emphasizes more human-like, proactive behaviors. Rather than passively waiting for user prompts, the model actively monitors the visual stream and generates timely outputs based on unfolding content. For instance, in Figure 1 (bottom), the assistant provides step-by-step guidance as the drawing progresses without being explicitly asked, simulating continuous support in dynamic environments.
To bridge the gap between offline and streaming video understanding, we introduce StreamBridge, a simple yet effective framework that seamlessly transforms pre-trained offline Video-LLMs into streaming-capable models. In contrast to prior efforts [10; 11; 12; 13], which train streaming models from scratch but fall behind on offline video tasks, StreamBridge leverages the strong generalization capabilities of existing Video-LLMs without requiring full retraining. Concretely, StreamBridge introduces a memory buffer to manage incoming video frames, coupled with a round-decayed compression strategy that merges earlier frame tokens while preserving recent ones, enabling the model to support long-context, multi-turn interactions in streaming scenarios. For proactive capabilities, instead of modifying the base model architecture [14] or introducing streaming-specific objectives [12], both of which can lead to optimization conflicts and issues like probability correction [10], StreamBridge adopts a modular design, by decoupling the proactive capability from the main Video-LLM via a compact activation model. This plug-and-play component operates in parallel with the main Video-LLM, enabling proactive behavior in a flexible and non-intrusive manner while fully preserving the main Video-LLM’s language fluency and general video understanding capabilities.
To further support StreamBridge, we construct Stream-IT, a large-scale dataset tailored for streaming scenarios. Stream-IT captures diverse real-time questions and proactive responses embedded within multi-turn video interactions, featuring interleaved video-text sequences. While existing datasets primarily focus on single-turn question answering [15; 16] or short-form video captioning [17; 18; 19], Stream-IT fills a critical gap by enabling temporally extended, interactive video understanding. It is constructed by concatenating semantically related short clips from large-scale video-caption corpora, followed by the generation of multi-turn QA sequences that simulate realistic, time-sensitive user interactions. Moreover, Stream-IT incorporates a broad spectrum of task formats sourced from public datasets, thereby boosting task diversity and promoting model generalization in streaming settings.
By integrating our StreamBridge framework and fine-tuning on Stream-IT, we successfully convert several leading offline Video-LLMs, including LLaVA-OV [3], Oryx-1.5 [1], and Qwen2-VL [2], into streaming-capable assistants. Extensive experiments demonstrate that our models achieve state-of-the-art performance on streaming benchmarks such as OVO-Bench [20] and Streaming-Bench [21], outperforming even proprietary models like GPT-4o [22] and Gemini 1.5 Pro [23], while retaining or exceeding performance on conventional offline video understanding tasks [24; 25; 26; 27; 28; 29].
2 Related Work
Video Large Language Models. With the rapid advancement of Multimodal Large Language Models (MLLMs) [30; 3; 31; 32; 2], Video-LLMs [33; 34; 35; 36; 15; 37; 38] have gained increasing attention for general video understanding. Typically, these models comprise a visual encoder [39; 40; 1] for extracting frame-level representations, a modality projector (e.g., MLP [41] and Q-former [30]) to map visual features into the language space, and an LLM [42; 43] to generate contextual responses. While achieving strong results on standard video benchmarks [25; 29; 27], these models are inherently designed for static, offline settings where the entire video is pre-recorded and fully accessible at inference time. As a result, they struggle in streaming environments, where video frames arrive sequentially and require real-time, temporally coherent, or even proactive responses. Our work aims to bridge this gap by augmenting offline Video-LLMs with streaming capabilities.
Streaming Video Understanding. Typical tasks in streaming video understanding, such as action recognition [44; 45; 46; 47] and anticipation [48; 49], causally process video inputs using only past and current observations. Recent efforts [50; 12; 13; 51] focus on building Video-LLMs capable of real-time conversation, generating timely responses throughout a live video stream. VideoLLM-Online [10] and Flash-VStream [11] introduce specialized online objectives and memory architectures to handle sequential inputs. MMDuet [14] and ViSpeak [52] add dedicated heads to facilitate proactive response generation. To benchmark streaming video capabilities, several evaluation suites have been proposed, including StreamingBench [21], StreamBench [53], SVBench [54], OmniMMI [55], and OVO-Bench [20]. In contrast to previous approaches that retrain models or tightly couple proactive mechanisms within the backbone, our work leverages the strong generalization abilities of pre-trained offline Video-LLMs [1; 3; 2]. We propose an efficient adaptation framework, combined with a dedicated fine-tuning dataset, to endow these models with streaming capabilities. Furthermore, observing that embedding the activation function into the main model often lead to optimization conflicts and performance degradation [10; 14], we advocate a modular, decoupled design. Our method introduces a compact, plug-and-play activation model that enables proactive behaviors efficiently and non-intrusively.
3 Methodology
3.1 Preliminary Analysis
Streaming video understanding involves interleaved video-text inputs. From an input perspective, streaming scenarios can be broadly categorized into two representative formats:
Multi-turn dialogue with interleaved video-text. In this setting, the input sequence is in the form of ‘<> <> <>, <> <> <>, ’, where <>, <>, and <> denote the video clip, user query, and assistant answer in the -th round. Crucially, there is no delay between <> and <>, reflecting the need for immediate responses. This format closely resembles the live interaction in dynamic environments, as shown in Figure 1 (Top).
Proactive output. The assistant answers after watching an incoming video stream, often without an explicit user query at the response time. The input can be structured as ‘<> <> <> <> <> ’, where <> represents an initial prompt (e.g.,“Guide me through the task”), and the model must proactively determine when and how to respond based on the incoming video contents. This scenario requires the ability to continuously monitor evolving context and trigger responses at appropriate moments. Figure 1 (Bottom) is an example of proactive responses.
Recent benchmarks such as OVO-Bench [20] and Streaming-Bench [21] attempt to evaluate these capabilities by constructing multi-turn interleaved video-text dialogues. However, due to the limited input length and the lack of streaming support in current Video-LLMs, these benchmarks necessarily simplify the problem. Specifically, they segment a complete long video into multiple isolated clips aligned with each query timestamp. For a query <> at time , the visual input is restricted to the uniformly sampled frames under segment , and prior dialogue history is completely discarded. As a result, the multi-turn streaming scenario is reduced to a series of independent, single-turn offline tasks. To address these limitations, we propose StreamBridge, a general framework designed to introduce the actual streaming setup to existing offline Video-LLMs.
3.2 StreamBridge
As shown in Figure 2 and Algorithm 1, in addition to the frame encoder and the large language model , StreamBridge proposes three key components to enable streaming capabilities: (1) a memory buffer responsible for storing and retrieving frame tokens over time, (2) a round-decayed compression strategy that efficiently prunes redundant tokens from earlier rounds while preserving the most recent context, and (3) a compact activation model that enables proactive responses by making frame-level decisions on when to generate outputs.
3.2.1 Memory Buffer
In streaming scenarios where video frames arrive sequentially, we adopt a memory buffer to store both visual and textual embeddings. As illustrated in Figure 2, each incoming frame is independently encoded and appended to the buffer alongside any associated query embeddings. Conceptually, operates under a producer-consumer paradigm: the encoder functions as the producer, continuously generating frame-level features, while the language model serves as the consumer, retrieving the accumulated embeddings to generate a response upon receiving a user query. Formally, as detailed in Algorithm 1, at each time step , the incoming frame is first processed by , and the resulting embeddings are stored in the memory buffer (Algorithm 1, line 4). Upon the arrival of a user query and a positive activation decision , the buffer content, including both visual and textual embeddings, is flattened into a single sequence of input embeddings, which is then fed into for response generation (Algorithm 1, line 13-16). Once a response is produced, it is also appended to the memory buffer (Algorithm 1, line 17), enabling the model to preserve temporal continuity and maintain a complete history of multi-turn video-text interactions.
3.2.2 Round-Decayed Compression
Online scenarios often involve long, even infinite video streaming, which can lead to significant memory usage and inference latency. Therefore, we propose a round-decayed token compression strategy tailored for multi-turn streaming settings. Specifically, we pre-define a maximum allowable embedding length for the model input. Before each response generation, the model checks whether the current input embedding exceeds . If so, we apply a round-decayed token merging strategy: starting from the earliest dialogue rounds, visual tokens are progressively merged frame-by-frame, until the total length falls below . The merging is implemented via average pooling [56] over adjacent frame tokens. This strategy ensures that the most recent visual context is retained with minimal distortion, thus maintaining the precision of real-time responses while not fully discarding historical visual contexts. At the same time, it significantly improves memory efficiency and reduces inference overhead as in Figure 5. This process is encapsulated in the compression function in Algorithm 1 (line 15). The detailed pseudo codes can be found in Appendix G.
3.2.3 A Plug-and-play Activation Model
To enable proactive responses, we decouple the activation function from the main Video-LLM. Unlike prior methods that tightly integrate activation mechanisms into the LLM [10; 12; 14; 52], our framework avoids potential interference with the language modeling capacity of the main Video-LLM. Specifically, we propose a parallel pipeline, where a compact external MLLM (e.g., LLaVA-OV-0.5B [3]) is used as an independent activation model, denoted as . As shown in Algorithm 1 (line 9), upon receiving each new frame, the framework simultaneously forwards the current frame (along with the user query and optionally previous frames) to to determine whether a response should be generated. If the activation signal is positive, the buffered embeddings are sent to the LLM for decoding. This design ensures high flexibility and compatibility. Furthermore, in real-time deployment, the , the frame encoder , and the main can run concurrently in parallel threads, enabling more efficient inference.
To train the activation model (illustrated in Figure 3), we modify the architecture by replacing the standard language modeling head with a score head for binary classification, and introduce a learnable activation token <> which is appended to the visual embeddings of each frame. After processing through the final layer, we extract the latest frame’s activation token and feed its hidden representation into the score head to predict whether the model should respond at that time. During inference, only when the predicted score is greater than the activation threshold , the main Video-LLM can be triggered to give a response. Since performs only binary classification (i.e., to respond or not), we aggressively pool its visual tokens for efficiency. The input sequence to the model follows the format: ‘<> <> <> <> <> ’, where the question is prepended to the sequence, and visual frames and corresponding responses are interleaved. This design enables the model to learn temporal dependencies and identify appropriate response moments throughout the video stream.
For training data, we collect a diverse set of temporally annotated video datasets across multiple tasks, including dense video captioning [57; 58], sequential step recognition [59; 60], grounded video question answering [61; 62; 63], temporal video grounding [64], and temporal action detection [65; 66]. For each task, we design specific prompt templates and randomly select among them as during training (details in Appendix A). To supervise activation timing, we insert the response <> at the end of its corresponding annotated timestamp (during inference, the response <> is generated by the larger main Video-LLM). Besides, only the last of frames of each video segment are labeled as positive (i.e., response-worthy), while earlier frames are treated as negatives. is dynamically sampled between 0% and 50% for each training instance, simulating a variety of activation patterns and enhancing the model’s robustness to temporal variations.
4 Stream-IT Dataset
As analyzed in Section 3.1, streaming scenarios are primarily characterized by multi-turn real-time understanding and proactive responses. However, existing datasets and video sources fall short of fully supporting these requirements [67; 68]. To fill this gap and further enhance the streaming interaction capability of StreamBridge, we introduce Stream-IT—a video-text dataset specifically designed for streaming instruction tuning with an interleaved multi-turn dialogue format.
Datasets for Proactive Understanding. We collect a set of public datasets enriched with timestamp annotations, spanning a wide range of tasks including: (i) Dense Video Captioning [57; 58; 69]; (ii) Sequential Step Recognition [59; 60; 70]; (iii) Grounded VideoQA [61; 71; 72; 73]. All datasets are reformatted into a proactive-style interleaved format: ‘<> <> <>, <> <>, ’, where may be an open-ended query (e.g., “Who is the man going to find?”) or a goal-oriented instruction (e.g., “Show me all the steps for cooking.”). Unlike traditional single-turn datasets where a question is immediately followed by an answer [67; 68], our structure introduces a temporal delay between <> and <> through the inserted video segments <>, simulating proactive response scenarios.
StreamingQA-120K: Multi-Turn, Long-Form QA Construction. To further support long-context, multi-turn real-time understanding, we introduce StreamingQA-120K, a large-scale synthetic dataset constructed by composing long-form videos from existing short video clips. Labeling long-duration videos with dense multi-turn QA pairs is prohibitively expensive. To address this, we leverage short clips from large-scale video-caption datasets, including WebVid-10M [19], Panda-70M [18], and InternVid-10M [17]. We filter approximately 1.28 million clips using semantic similarity between video and caption to ensure alignment, with each clip being around 12 seconds long. With these short clips, to form coherent long-form videos, we then iteratively compute pairwise similarity between videos and concatenate highly similar clips. Each constructed video contains roughly 10 clips, with an average length exceeding 150 seconds. Captions for each clip are preserved with natural timestamps. Using these captions, we employ GPT-4o [22] to generate diverse question-answer pairs spanning 8 task types. By default, each QA pair <> <> is inserted immediately after its corresponding clip <>, forming sequences like ‘<> <> <>, <> <> <>, …’. Here, we introduce two augmentation strategies during sequence construction:
Random QA Drop: randomly drops some QA pairs by transforming ‘<> <> <>’ to ‘<>’ with a probability of , to prevent overfitting to fixed QA positions and enhance the model’s robustness in temporal variations. We set to be 0.55 by default.
QA Interval Shift: with probability , transforms sequences from ‘<> <> <>’ to ‘<> <> <>’, where the visual content serves as the temporal delay between question and response for proactive scenarios. is set to 0.1 here.
Together, these strategies ensure that the Stream-IT dataset supports rich and varied streaming interaction formats, enabling both multi-turn real-time dialogue and proactive response capabilities across a wide range of tasks and timescales. More details on data statistics, concatenation strategy of StreamingQA-120K, and prompts for QA generation are provided in Appendix B.
| Method | # of | OVO-Bench Real-Time | Streaming-Bench Real-Time | ||||||||||||||||
| Frames | OCR | ACR | ATR | STU | FPD | OJR | AVG. | OP | CR | CS | ATP | EU | TR | PR | SU | ACP | CT | AVG. | |
| Human | |||||||||||||||||||
| Human | - | 93.96 | 92.57 | 94.83 | 92.70 | 91.09 | 94.02 | 93.20 | 89.47 | 92.00 | 93.60 | 91.47 | 95.65 | 92.52 | 88.00 | 88.75 | 89.74 | 91.30 | 91.46 |
| Proprietary Models (Offline), Single-Turn Evaluation | |||||||||||||||||||
| Gemini 1.5 pro [23] | 1 FPS | 85.91 | 66.97 | 79.31 | 58.43 | 63.37 | 61.96 | 69.32 | 79.02 | 80.47 | 83.54 | 79.67 | 80.00 | 84.74 | 77.78 | 64.23 | 71.95 | 48.70 | 75.69 |
| GPT-4o [22] | 64 | 69.80 | 64.22 | 71.55 | 51.12 | 70.3 | 59.78 | 64.46 | 77.11 | 80.47 | 83.91 | 76.47 | 70.19 | 83.80 | 66.67 | 62.19 | 69.12 | 49.22 | 73.28 |
| Open-Source Models (Offline), Single-Turn Evaluation | |||||||||||||||||||
| Qwen2-VL-72B [2] | 64 | 65.77 | 60.55 | 69.83 | 51.69 | 69.31 | 54.35 | 61.92 | - | - | - | - | - | - | - | - | - | - | - |
| LLaVA-Video-7B [15] | 64 | 69.13 | 58.72 | 68.83 | 49.44 | 74.26 | 59.78 | 63.52 | - | - | - | - | - | - | - | - | - | - | - |
| LLaVA-OV-7B [3] | 64/32 | 66.44 | 57.80 | 73.28 | 53.37 | 71.29 | 61.96 | 64.02 | 80.38 | 74.22 | 76.03 | 80.72 | 72.67 | 71.65 | 67.59 | 65.45 | 65.72 | 45.08 | 71.12 |
| Qwen2-VL-7B [2] | 64/1 FPS | 60.40 | 50.46 | 56.03 | 47.19 | 66.34 | 55.43 | 55.98 | 75.20 | 82.81 | 73.19 | 77.45 | 68.32 | 71.03 | 72.22 | 61.19 | 61.47 | 46.11 | 69.04 |
| InternVL-V2-8B [74] | 64/16 | 67.11 | 60.55 | 63.79 | 46.07 | 68.32 | 56.52 | 60.39 | 68.12 | 60.94 | 69.40 | 77.12 | 67.70 | 62.93 | 59.26 | 53.25 | 54.96 | 56.48 | 63.72 |
| Open-Source Models (Streaming), Single-Turn Evaluation | |||||||||||||||||||
| Flash-VStream-7B [11] | 1 FPS | 24.16 | 29.36 | 28.45 | 33.71 | 25.74 | 28.80 | 28.37 | 25.89 | 43.57 | 24.91 | 23.87 | 27.33 | 13.08 | 18.52 | 25.20 | 23.87 | 48.70 | 23.23 |
| VideoLLM-Online-8B [10] | 2 FPS | 8.05 | 23.85 | 12.07 | 14.04 | 45.54 | 21.20 | 20.79 | 39.07 | 40.06 | 34.49 | 31.05 | 45.96 | 32.40 | 31.48 | 34.16 | 42.49 | 27.89 | 35.99 |
| Dispider [13] | 1 FPS | 57.72 | 49.54 | 62.07 | 44.94 | 61.39 | 51.63 | 54.55 | 74.92 | 75.53 | 74.10 | 73.08 | 74.44 | 59.92 | 76.14 | 62.91 | 62.16 | 45.80 | 67.63 |
| Models under StreamBridge (Offline Streaming), Multi-Turn Evaluation | |||||||||||||||||||
| Oryx-1.5-7B† [1] | 1 FPS | 60.40 | 52.29 | 69.83 | 50.00 | 65.35 | 57.61 | 59.25 | 78.47 | 77.17 | 83.86 | 80.20 | 71.07 | 66.98 | 79.63 | 61.38 | 66.29 | 40.93 | 70.59 |
| + Stream-IT | 1 FPS | 84.56 | 75.23 | 70.69 | 50.56 | 74.26 | 71.74 | 71.17 | 82.29 | 77.95 | 87.98 | 86.47 | 77.99 | 81.31 | 76.85 | 69.92 | 71.96 | 35.23 | 74.79 |
| LLaVA-OV-7B† [3] | 1 FPS | 58.39 | 59.63 | 69.82 | 44.38 | 76.23 | 61.41 | 61.64 | 76.84 | 77.17 | 82.60 | 75.25 | 64.15 | 64.17 | 75.00 | 61.38 | 61.19 | 46.11 | 68.39 |
| + Stream-IT | 1 FPS | 74.50 | 77.06 | 70.69 | 54.49 | 73.27 | 69.57 | 69.93 | 82.29 | 72.44 | 92.09 | 80.86 | 71.07 | 74.46 | 75.00 | 62.20 | 70.26 | 28.50 | 70.92 |
| Qwen2-VL-7B† [2] | 1 FPS | 65.10 | 64.22 | 64.66 | 46.63 | 74.26 | 65.22 | 63.35 | 80.38 | 78.74 | 83.22 | 79.86 | 74.21 | 69.47 | 77.78 | 63.41 | 69.97 | 43.01 | 72.01 |
| + Stream-IT | 1 FPS | 84.56 | 71.56 | 74.14 | 49.44 | 75.25 | 72.83 | 71.30 | 84.74 | 82.68 | 88.92 | 89.77 | 77.36 | 85.36 | 84.26 | 69.92 | 71.67 | 35.75 | 77.04 |
5 Experiments
5.1 Settings
Models and Datasets. We evaluate StreamBridge framework using three mainstream offline Video-LLMs to show its generalizability: LLaVA-OV-7B [3], Qwen2-VL-7B [2], and Oryx-1.5-7B [1]. To preserve their general video understanding capabilities during streaming adaptation, we supplement Stream-IT with approximately 600K samples from the LLaVA-178K [15], VCG-Plus [35] and ShareGPT4Video [16]. For the activation model, we fine-tune LLaVA-OV-0.5B [3] on our collected activation datasets as described in Sec. 3.2.3. The videos are sampled at 1 FPS. In Section 5.3, we use Qwen2-VL-7B as the default model unless otherwise specified. See the Appendix C for more details.
Benchmarks. For multi-turn real-time understanding, we choose OVO-Bench [20] and Streaming-Bench [21]. We primarily focus on their real-time tasks. For general video understanding, we evaluate our method across 7 video benchmarks, including 3 short-video benchmarks: MVBench [24], PerceptionTest [26], TempCompass [75], and 4 long-video benchmarks: EgoSchema [28], LongVideoBench [29], MLVU [27], and VideoMME [25]. To evaluate the proactive capability of our method, we use subtasks from ET-Bench [64] following previous works. See Appendix D for more benchmark details and evaluation metrics.
5.2 Main Results
Multi-Turn Real-Time Understanding. As discussed in Section 3.1, the results reported in the original paper [20; 21] in Table 1, marked as “(Offline), Single-Turn Evaluation”, do not reflect performance in real streaming scenarios. They segment a complete video into several individual clips, discarding historical visual and dialogue contexts, thereby limiting the upper bound of the performance. In contrast, with the StreamBridge framework, denoted as “(Offline Streaming), Multi-Turn Evaluation”, these offline models are equipped to process streaming videos at 1 FPS in a multi-turn manner, while maintaining input length and historical contexts within a predefined maximum token budget.
Specifically, we observe that Qwen2-VL† demonstrates notable improvements in the streaming setting, with its average score on OVO-Bench increasing from 55.98 to 63.35, and on Streaming-Bench from 69.04 to 72.01. Conversely, LLaVA-OV† shows a slight performance drop when transitioning to the streaming setup: from 64.02 to 61.64 on OVO-Bench, and from 71.12 to 68.39 on Streaming-Bench. We attribute these differences to the nature of their pretraining data, where Qwen2-VL benefits from richer interleaved multimodal training (e.g., image/video-text sequences), which makes it more adept at understanding interleaved video-text inputs and utilizing extended context effectively. On the other hand, LLaVA-OV is trained with fewer interleaved sequences, making it less suited for multi-turn streaming inputs. When faced with long, interleaved video-text sequences in streaming scenarios, its performance tends to degrade as more historical frames accumulate. Notably, fine-tuning these models on the proposed Stream-IT leads to substantial improvements in multi-turn real-time understanding. For instance, Oryx-1.5† achieves a performance gain of +11.92 on OVO-Bench and +4.2 on Streaming-Bench. Furthermore, Qwen2-VL† reaches an average score of 71.30 on OVO-Bench and 77.04 on Streaming-Bench, outperforming proprietary models such as GPT-4o and Gemini 1.5 Pro. These results validate the effectiveness of both our StreamBridge framework and the Stream-IT dataset in enhancing multi-turn real-time understanding in streaming scenarios.
| Model | MVBench | PerceptionTest | TempCompass | EgoSchema | LongVideoBench | MLVU | VideoMME (w/o subs) |
| Avg | Val | MC | Test | Val | M-Avg | Avg | |
| Avg. Duration | 16s | 23s | 12s | 180s | 473s | 651s | 1010s |
| Proprietary Models | |||||||
| Gemini 1.5 pro [23] | 60.5 | - | 67.1 | 71.2 | 64.0 | - | 75.0 |
| GPT-4o [22] | 64.6 | - | 70.9 | 72.2 | 66.7 | 64.6 | 71.9 |
| Open-Source Models | |||||||
| Kangaroo-8B [76] | 61.0 | - | 62.5 | - | 54.8 | 61.0 | 56.0 |
| LongVILA-7B [77] | - | - | - | 67.7 | - | - | 57.5 |
| LongVU-7B [78] | 66.9 | - | - | 67.6 | - | 65.4 | 60.6 |
| Apollo-7B [4] | - | 67.3 | 64.9 | - | 58.5 | 70.9 | 61.3 |
| NVILA-8B [79] | 68.1 | 65.4 | 69.7 | - | 57.7 | 70.1 | 64.2 |
| SF-LLaVA-1.5-7B [5] | - | 69.6 | 68.8 | - | 62.5 | 71.5 | 63.9 |
| InternVL2.5-8B [80] | 72.0 | 68.2 | 68.3 | 51.5 | 60.0 | 68.9 | 64.2 |
| VideoChat-Flash-7B [81] | 74.0 | 76.2 | - | - | 64.7 | 74.7 | 65.3 |
| VideoLLaMA3-7B [37] | 69.7 | 72.8 | 68.1 | 63.3 | 59.8 | 73.0 | 66.2 |
| Oryx-1.5-7B [1] | 67.6 | 70.0 | 58.8 | - | 56.3 | 67.5 | 58.8 |
| Oryx-1.5-7B (ours) ‡ | 68.0 (0.4) | 71.0 (1.0) | 69.0 (10.2) | 61.2 | 58.9 (2.6) | 71.4 (4.0) | 65.5 (6.7) |
| LLaVA-OV-7B [3] | 56.7 | 57.1 | 64.8 | 60.1 | 56.3 | 64.7 | 58.2 |
| LLaVA-OV-7B (ours) ‡ | 59.4 (2.7) | 63.9 (6.8) | 67.7 (2.9) | 67.0 (6.9) | 54.3 (2.0) | 68.2 (3.5) | 61.2 (3.0) |
| Qwen2-VL-7B [2] | 67.0 | 62.3 | 67.9 | 66.7 | - | - | 63.3 |
| Qwen2-VL-7B (ours) ‡ | 64.4 (2.6) | 69.9 (7.6) | 71.1 (3.2) | 66.9 (0.2) | 59.1 | 69.6 | 64.4 (1.1) |
General Video Understanding. While our StreamBridge is designed for online scenarios, we also verify that it does not downgrade the base model’s performance on standard offline video tasks. As shown in Table 2, models equipped with the StreamBridge framework and fine-tuned on Stream-IT (denoted with ‡) exhibit consistent improvements or maintain comparable performance relative to their original versions. For instance, Oryx-1.5-7B‡ achieves 65.5 on the challenging VideoMME with an increase of 6.7, while LLaVA-OV-7B‡ outperforms its base model across nearly all benchmarks, except LongVideoBench. Likewise, Qwen2-VL-7B‡ achieves competitive results on MVBench, while surpassing its original counterpart on other benchmarks. These outcomes demonstrate that our streaming adaptation enables models to retain, or even exceed their original performance in general video understanding tasks, demonstrating the generality and non-degradability of our method.
| Method | # of | ET-Bench | |||||
| Frames | TVGF1 | TALF1 | DVCF1 | DVCSim | SLCF1 | SLCSim | |
| VideoLLM-Online [10] | 2 FPS | 13.2 | 9.1 | 24.0 | 13.4 | 9.9 | 10.1 |
| Dispider [13] | 1 FPS | 36.1 | 27.3 | 33.8 | 18.9 | 18.8 | 12.4 |
| Models under StreamBridge Framework | |||||||
| Oryx-1.5 (ours)‡ | 1 FPS | 34.3 | 24.3 | 37.8 | 24.0 | 22.5 | 17.3 |
| LLaVA-OV (ours)‡ | 1 FPS | 34.3 | 24.3 | 37.9 | 24.2 | 22.8 | 16.2 |
| Qwen2-VL (ours)‡ | 1 FPS | 34.3 | 24.3 | 38.3 | 25.1 | 22.6 | 17.1 |
Online Activation. We evaluate the proactive capability of our framework in Table 3. Notably, in all tasks, the question is presented at the beginning of the video, and the model must autonomously decide when to respond. On the ET-Bench, StreamBridge outperforms both VideoLLM-Online [10] and Dispider [13] across generation-based tasks such as DVC (Dense Video Captioning) and SLC (Step Localization and Captioning), achieving higher similarity scores of DVCSim and SLCSim, by producing more accurate and context-aware descriptions in streaming scenarios. We attribute this to the decoupled nature of the activation model, which enables the main Video-LLM to focus solely on video understanding and language generation, free from the burden of proactive decision-making. We also observe that Qwen2-VL‡ achieves better text similarity scores than other Video-LLMs, consistent with its strong real-time understanding performance presented in Table 1.
5.3 In-Depth Analysis
| Compression | OVO | Streaming | ET | |
| Avg. | Avg. | DVCSim | SLCSim | |
| Truncation | 68.88 | 72.79 | 22.1 | 16.7 |
| Round-Uniform | 69.91 | 74.18 | 23.8 | 15.9 |
| Round-Decayed | 71.30 | 77.04 | 25.1 | 17.1 |
| LLaVA-178k | Stream-IT | OVO | Streaming | MVBench | VideoMME | |
| (600k used) | w/o SQA-120k | w/ SQA-120k | Avg. | Avg. | Avg. | Overall. |
| 65.98 | 71.36 | 64.5 | 61.7 | |||
| 71.28 | 74.10 | 58.8 | 59.0 | |||
| 67.67 | 72.42 | 63.1 | 63.6 | |||
| 71.30 | 77.04 | 64.4 | 64.4 | |||
Round-Decayed Compression. We set the maximum input length , and denote the current length of the input embeddings as . To assess the effectiveness of our round-decayed compression strategy, we compare it against two alternative methods: (1) Truncation: If , only keep the last tokens in the input sequence. (2) Round-Uniform: We treat each round equally by reducing the number of visual tokens with a fixed ratio per round, to keep the total length within . The results are reported in Table 5. We observe that Truncation yields the worst performance, as it indiscriminately removes both visual and textual history tokens, severely weakening multi-turn reasoning. The Round-Uniform strategy performs slightly better, but still underperforms our method. It compresses the latest visual tokens, which are critical for real-time comprehension, thus leading to degraded performance, particularly on OVO-Bench and Streaming-Bench.
Inference Latency. We also evaluate the inference latency on a single A100-80G GPU with different (8k, 16k, 32k), as shown in Figure 5. Our results show that our compression method maintains near-constant latency when the number of input tokens exceeds , whereas models without compression suffer from sharply increasing delays and eventually trigger out-of-memory (OOM) errors with 2048 frames. This highlights the necessity of effective compression to balance inference efficiency and memory usage in streaming settings.
Impact of Stream-IT. Table 5 ablates effectiveness of Stream-IT. Training on LLaVA-178K alone causes a marked drop on both OVO-Bench and Streaming-Bench, as it lacks interleaved video–text samples necessary for multi-turn interactions. Conversely, using only Stream-IT without LLaVA-178K leads to declines in general video understanding (MVBench, VideoMME), indicating that the larger offline data corpus still contributes valuable world knowledge. Finally, removing the synthetic StreamingQA-120K subset from Stream-IT degrades performance across both streaming and offline benchmarks, underscoring the crucial role of StreamingQA-120K in boosting both streaming and offline video understanding capabilities.
Activation Threshold. The compact activation model makes a per-frame decision to trigger responses, with frequency determined by the activation threshold (see score head in Figure 3). We adopt a default of 0.35, following common practice [52; 14]. Figure 5 illustrates the impact of varying this threshold: both excessively low and high values of decrease F1 scores ( and on ET-Bench). A low threshold triggers overly frequent responses, while a high threshold suppresses them excessively, both of which hurt performance. Nonetheless, this hyper-parameter allows users to flexibly control response frequency through , adapting to different practical scenarios.
6 Conclusion
We present StreamBridge, a novel framework that transforms offline Video-LLMs into streaming-capable models. StreamBridge introduces a memory buffer paired with a round-decayed compression strategy, and decouples the activation function with a compact activation model. We also construct Stream-IT, a dataset with interleaved video-text sequences to further support StreamBridge. Extensive experiments on diverse benchmarks demonstrate that our method not only preserves the strengths of the base models but also equips them with the ability to make timely, proactive responses across multi-turn, long-context streaming scenarios. We believe StreamBridge offers a general solution for bridging the gap between offline Video-LLMs and real-world, interactive streaming applications.
Appendix A Datasets Used to Train the Activation Model
To train the activation model , we compile a diverse collection of video datasets spanning five distinct tasks:
In total, our training set contains approximately 180k video samples. For each sample, we construct an input prompt using task-specific templates. A prompt is randomly sampled from a predefined pool for the corresponding task to ensure stylistic diversity and improve generalization across video domains. The full list of prompt templates is provided in Table 6. During training, the prompt is inserted at the beginning of the input sequence as in Figure 3.
Appendix B Stream-IT Construction
B.1 Statistics of Stream-IT
We provide detailed statistics of the Stream-IT dataset in Table 8, including the number of samples, average video duration, and the corresponding source datasets used for each task. Notably, during the construction of the dense video captioning tasks, including ActivityNet[57] and Shot2Story [58], we only arrange 20% of the sequences with the proactive format of ‘<> <> <>, <> <>, ‘, while the 80% of the sequences with the multi-turn format of ’<> <> <>, <> <> <>, …’, where <> is the question asking about current situations like ’What is happening now?’.
B.2 Concatenation Strategy for Constructing StreamingQA-120K
Starting from a pool of 1.28 million filtered short videos sourced from WebVid-10M [19], Panda-70M [18], and InternVid-10M [17], our goal is to iteratively merge semantically similar clips to form long-form video samples. Let denote the entire set of filtered videos. We initiate the process by randomly sampling one clip from as the anchor. We then compute pairwise semantic similarity (based on the middle frame) between and all other videos in . A new clip is sampled according to the similarity distribution (without replacement). The procedure is repeated using as the new anchor, generating from , and so on. This results in a similarity-ordered list of videos . This approach allows for flexible concatenation of any number of clips to construct a long-form sample, by directly selecting a continuous span , without re-computing similarity each time. We also prepare hallucination questions irrelevant to existing video inputs following [20] with a ratio of 0.01%.
B.3 Prompt Templates for Generating QA Pairs
To generate question-answer pairs based on clip-level captions, we design diverse prompt templates for 8 distinct reasoning tasks. Table 7 provides examples of these templates. Below, we summarize the <task descriptions> associated with each QA category:
[OP] Object Perception: Detect and identify objects, focusing on recognizing their attributes in real time.
[AR] Action Recognition: Identify human actions and interactions occurring in the current moment.
[SA] Spatial Awareness: Understand spatial relationships among objects and events; reason about location, orientation, and distance.
[SR] Sequential Relationship: Identify the temporal order of events and actions, especially those involving “before” and “after” cues.
[CR] Causal Reasoning: Analyze cause-and-effect relationships between actions and outcomes.
[OCR] Optical Character Recognition: Recognize and interpret textual content in scenes (e.g., subtitles, signs, charts).
[UEH] Unexpected Event Handling: Detect and react to anomalies or sudden changes in the environment.
[EU] Event Understanding: Summarize and reason about sequences of temporally linked events.
These diverse prompts ensure broad task coverage and help enhance the model’s generalization across different temporal and semantic understanding challenges.
| Task | # of Samples | Datasets | Average duration |
| Dense Video Captioning | 54k | ActivityNet [57] (10k) | 180s |
| Shot2Story [58] (36k) | 16s | ||
| ViTT [69] (8k) | 210s | ||
| Sequential Step Recognition | 22k | YouCook2 [60] (1.3k) | 317s |
| COIN [59] (11k) | 145s | ||
| HowToStep [70] (10k) | 190s | ||
| Grounded Video Question Answering | 69k | MovieChat [72] (0.8k) | 10k frames |
| EgoTimeQA [71] (10k) | 150s | ||
| QAEgo4D [61] (15k) | 495s | ||
| FineVideo [73] (43k) | 280s | ||
| Multi-turn Real-time Question Answering | 120k | StreamingQA-120K (120k) | 150s |
| (Sourced from Webvid-10M[19], Panda-70M[18], InternVid-10M[17]) |
Appendix C More Implementation Details
For the main VideoLLMs, we use the following configurations for each model:
LLaVA-OV-7B: We apply center cropping with a resolution of 384384 and use a 4 down sampler (bilinear interpolation) with the frame features, resulting in 49 tokens per frame.
Oryx-1.5-7B: We use the model’s default dynamic resolution, ranging from 288 to 480 pixels. With a 4 down sampler on the frame features (average pooling), the resulting token count per frame varies between 33 and 59.
Qwen2-VL-7B: The model uses a dynamic resolution between 224 and 448, with 4 down sampling (average pooling) on the frame features, resulting in 36–64 tokens per frame.
All models are fine-tuned for one epoch using a learning rate of 2e-5 with a cosine annealing scheduler and AdamW optimizer. The image encoder remains frozen, while the visual projector and the LLM are fully trainable. The maximum length of input embeddings is set to 16384 for the round-decayed compression.
For the activation model, we adopt LLaVA-OV-0.5B as the base model. To improve efficiency, we aggressively pool the frame representations to 16 tokens per frame. During training, only the LoRA adapters, the projector, the score head, and the learnable activation token are trainable. The model is trained for 5 epochs using a fixed learning rate of 2e-5 for the projector, while 2e-4 for the LoRA adapters, score head, and the learnable activation token, with AdamW optimizer. Notably, for both the main VideoLLM and the activation model, we sample frames at 1 FPS to better simulate real-world frame rates. For videos longer than 256 seconds, we uniformly sample 256 frames to fit within the maximum input length constraint. Experiments are conducted on NVIDIA-H100/A100 GPUs. During inference, we sample videos at 2 FPS for short video benchmarks like MVBench, PerpecptionTest, and TempCompass, while 1 FPS for multi-turn real-time understanding benchmarks and long video benchmarks including OVO-Bench, Streaming-Bench, MLVU, LongVideoBench, VideoMME, and EgoSchema.
Appendix D Benchmarks and Metrics
Multi-turn Real-time Understanding Benchmarks. We evaluate our method on two recently proposed large-scale streaming video benchmarks: OVO-Bench [20] and Streaming-Bench [21]. Both benchmarks are designed to assess streaming video comprehension under long-context, multi-turn settings. Our evaluation primarily focuses on their real-time understanding tasks. OVO-Bench contains 512 videos with an average length of 435 seconds and approximately 1,600 questions. The evaluated tasks include: (1) Spatial Understanding (STU), (2) Object Recognition (OJR), (3) Attribute Recognition (ATR), (4) Action Recognition (ACR), (5) Optical Character Recognition (OCR), and (6) Future Prediction (FPD). Streaming-Bench consists of 500 videos with an average length of 606 seconds and approximately 2,500 questions. It includes the following tasks: (1) Object Perception (OP), (2) Causal Reasoning (CR), (3) Clip Summarization (CS), (4) Attribute Perception (ATP), (5) Event Understanding (EU), (6) Text-Rich Understanding (TR), (7) Prospective Reasoning (PR), (8) Spatial Understanding (SU), (9) Action Perception (ACP), (10) Counting (CT). Both benchmarks are structured as multiple-choice question-answering tasks, and we report the accuracy.
General Video Understanding Benchmarks. To evaluate general video comprehension ability, we test our models on seven widely used benchmarks. This includes three short-video benchmarks: MVBench [24], Perception Test [26], and TempCompass [75], and four long-video benchmarks: EgoSchema [28], LongVideoBench [29], MLVU [27], and VideoMME [25]. These datasets span a broad range of video durations, from a few minutes to several hours. All are evaluated in a multiple-choice format, and accuracy is reported.
Online Activation Benchmarks. To assess the proactive capabilities of our framework, we evaluate performance on a subset of ET-Bench [64], including Temporal Video Grounding (TVG), Temporal Action Localization (TAL), Dense Video Captioning (DVC), and Sequential Localization Captioning (SLC). These tasks emphasize a shift from passive to active perception, requiring the model to determine when to respond based on upcoming visual inputs, rather than reacting immediately. For example, the Sequential Localization Captioning (SLC) task requires the model to both determine the precise timing of a certain step and output its content. For evaluation metrics, we compute the average F1 score across multiple IoU thresholds (IoU {}) for localization-based tasks. For tasks involving text generation, we adopt sentence-level similarity metrics [83] to measure the semantic alignment between model outputs and ground-truth responses, following prior works [64; 13]. Specifically, the all-MiniLM-L6-v2 model in Sentence-Transformers library is used as the embedding model. Notably, in all these tasks, the question is presented at the beginning of the video, and the model must autonomously decide when to respond. Moreover, the results of TVGF1, TALF1, are the same for our method with different main Video-LLMs, since they use the same activation model and will not be affected by the generated response.
Appendix E Full Results on OVO-Bench and Streaming-Bench
| Method | # of | Real-Time Visual Perception | Backward Tracing | Forward Active Responding | Overall. | ||||||||||||
| Frames | OCR | ACR | ATR | STU | FPD | OJR | AVG. | EPM | ASI | HLD | AVG. | REC | SSR | CRR | AVG. | Overall AVG. | |
| Human | |||||||||||||||||
| Human | - | 93.96 | 92.57 | 94.83 | 92.70 | 91.09 | 94.02 | 93.20 | 92.59 | 93.02 | 91.37 | 92.33 | 95.48 | 89.67 | 93.56 | 92.90 | 92.81 |
| Proprietary Models (Offline), Single-Turn Evaluation | |||||||||||||||||
| Gemini 1.5 pro [23] | 1 FPS | 85.91 | 66.97 | 79.31 | 58.43 | 63.37 | 61.96 | 69.32 | 58.59 | 76.35 | 52.64 | 62.54 | 35.53 | 74.24 | 61.67 | 57.15 | 63.00 |
| GPT-4o [22] | 64 | 69.80 | 64.22 | 71.55 | 51.12 | 70.30 | 59.78 | 64.46 | 57.91 | 75.68 | 48.66 | 60.75 | 27.58 | 73.21 | 59.40 | 53.40 | 59.54 |
| Open-Source Models (Offline), Single-Turn Evaluation | |||||||||||||||||
| Qwen2-VL-72B [2] | 64 | 65.77 | 60.55 | 69.83 | 51.69 | 69.31 | 54.35 | 61.92 | 52.53 | 60.81 | 57.53 | 56.95 | 38.83 | 64.07 | 45.00 | 49.30 | 56.27 |
| LLaVA-Video-7B [15] | 64 | 69.13 | 58.72 | 68.83 | 49.44 | 74.26 | 59.78 | 63.52 | 56.23 | 57.43 | 7.53 | 40.4 | 34.10 | 69.95 | 60.42 | 54.82 | 52.91 |
| LLaVA-OV-7B [3] | 64 | 66.44 | 57.80 | 73.28 | 53.37 | 71.29 | 61.96 | 64.02 | 54.21 | 55.41 | 21.51 | 43.71 | 25.64 | 67.09 | 58.75 | 50.50 | 52.74 |
| Qwen2-VL-7B [2] | 64 | 60.40 | 50.46 | 56.03 | 47.19 | 66.34 | 55.43 | 55.98 | 47.81 | 35.48 | 56.08 | 46.46 | 31.66 | 65.82 | 48.75 | 48.74 | 50.39 |
| InternVL-V2-8B [74] | 64 | 67.11 | 60.55 | 63.79 | 46.07 | 68.32 | 56.52 | 60.39 | 48.15 | 57.43 | 24.73 | 43.44 | 26.5 | 59.14 | 54.14 | 46.60 | 50.15 |
| Open-Source Models (Streaming), Single-Turn Evaluation | |||||||||||||||||
| Flash-VStream-7B [11] | 1 FPS | 24.16 | 29.36 | 28.45 | 33.71 | 25.74 | 28.80 | 28.37 | 39.06 | 37.16 | 5.91 | 27.38 | 8.02 | 67.25 | 60.00 | 45.09 | 33.61 |
| VideoLLM-Online-8B [10] | 2 FPS | 8.05 | 23.85 | 12.07 | 14.04 | 45.54 | 21.20 | 20.79 | 22.22 | 18.80 | 12.18 | 17.73 | - | - | - | - | - |
| Dispider [13] | 1 FPS | 57.72 | 49.54 | 62.07 | 44.94 | 61.39 | 51.63 | 54.55 | 48.48 | 55.41 | 4.30 | 36.06 | 18.05 | 37.36 | 48.75 | 34.72 | 41.78 |
| Models under StreamBridge (Offline Streaming), Multi-Turn Evaluation | |||||||||||||||||
| Oryx-1.5-7B† [1] | 1 FPS | 60.40 | 52.29 | 69.83 | 50.00 | 65.35 | 57.61 | 59.25 | 54.21 | 55.41 | 5.40 | 38.33 | 20.65 | 37.56 | 40.00 | 32.74 | 43.44 |
| + Stream-IT | 1 FPS | 84.56 | 75.23 | 70.69 | 50.56 | 74.26 | 71.74 | 71.17 | 69.02 | 59.50 | 79.03 | 69.17 | 20.51 | 66.89 | 60.41 | 49.27 | 63.21 |
| LLaVA-OV-7B† [3] | 1 FPS | 58.39 | 59.63 | 69.82 | 44.38 | 76.23 | 61.41 | 61.64 | 53.87 | 54.72 | 30.64 | 46.41 | 14.41 | 51.23 | 43.33 | 36.33 | 48.13 |
| + Stream-IT | 1 FPS | 74.50 | 77.06 | 70.69 | 54.49 | 73.27 | 69.57 | 69.93 | 66.67 | 6149 | 85.48 | 71.21 | 17.83 | 66.06 | 61.67 | 48.52 | 63.22 |
| Qwen2-VL-7B† [2] | 1 FPS | 65.10 | 64.22 | 64.66 | 46.63 | 74.26 | 65.22 | 63.35 | 55.56 | 60.14 | 62.90 | 59.53 | 22.14 | 61.12 | 49.58 | 44.28 | 55.72 |
| + Stream-IT | 1 FPS | 84.56 | 71.56 | 74.14 | 49.44 | 75.25 | 72.83 | 71.30 | 67.68 | 57.43 | 79.03 | 68.05 | 19.17 | 64.25 | 61.67 | 48.36 | 62.57 |
| Method | # of | Real-Time Visual Understanding | Omni-Source Understanding | Contextual Understanding | Overall. | ||||||||||||||||||
| Frames | OP | CR | CS | ATP | EU | TR | PR | SU | ACP | CT | AVG. | ER | SCU | SD | MA | AVG. | ACU | MCU | SQA | PO | AVG. | Overall AVG. | |
| Human | |||||||||||||||||||||||
| Human | - | 89.47 | 92.00 | 93.60 | 91.47 | 95.65 | 92.52 | 88.00 | 88.75 | 89.74 | 91.30 | 91.46 | 88.00 | 88.24 | 93.60 | 90.27 | 90.26 | 88.80 | 90.40 | 95.00 | 100 | 93.55 | 91.66 |
| Proprietary Models (Offline), Single-Turn Evaluation | |||||||||||||||||||||||
| Gemini 1.5 pro [23] | 1 FPS | 79.02 | 80.47 | 83.54 | 79.67 | 80.00 | 84.74 | 77.78 | 64.23 | 71.95 | 48.70 | 75.69 | 46.80 | 39.60 | 74.90 | 80.00 | 60.22 | 51.41 | 40.73 | 54.80 | 45.10 | 48.73 | 67.07 |
| GPT-4o [22] | 64 | 77.11 | 80.47 | 83.91 | 76.47 | 70.19 | 83.80 | 66.67 | 62.19 | 69.12 | 49.22 | 73.28 | 41.20 | 37.20 | 43.60 | 56.00 | 44.50 | 41.20 | 38.40 | 32.80 | 56.86 | 38.70 | 60.15 |
| Open-Source Models (Offline), Single-Turn Evaluation | |||||||||||||||||||||||
| LLaVA-OV-7B [3] | 32 | 80.38 | 74.22 | 76.03 | 80.72 | 72.67 | 71.65 | 67.59 | 65.45 | 65.72 | 45.08 | 71.12 | 40.80 | 37.20 | 33.60 | 44.80 | 38.40 | 35.60 | 36.00 | 27.27 | 29.55 | 32.74 | 56.36 |
| Qwen2-VL-7B [2] | 0.2-1 FPS | 75.20 | 82.81 | 73.19 | 77.45 | 68.32 | 71.03 | 72.22 | 61.19 | 61.47 | 46.11 | 69.04 | 41.20 | 22.00 | 32.80 | 43.60 | 34.90 | 31.20 | 26.00 | 39.60 | 22.73 | 31.66 | 54.14 |
| InternVL-V2-8B [74] | 16 | 68.12 | 60.94 | 69.40 | 77.12 | 67.70 | 62.93 | 59.26 | 53.25 | 54.96 | 56.48 | 63.72 | 37.60 | 26.40 | 37.20 | 42.00 | 35.80 | 32.00 | 31.20 | 32.32 | 40.91 | 32.42 | 51.40 |
| Open-Source Models (Streaming), Single-Turn Evaluation | |||||||||||||||||||||||
| Flash-VStream-7B [11] | 1 FPS | 25.89 | 43.57 | 24.91 | 23.87 | 27.33 | 13.08 | 18.52 | 25.20 | 23.87 | 48.70 | 23.23 | 25.91 | 24.90 | 25.60 | 28.40 | 26.00 | 24.80 | 25.20 | 26.80 | 1.96 | 24.12 | 24.04 |
| VideoLLM-Online-8B [10] | 2 FPS | 39.07 | 40.06 | 34.49 | 31.05 | 45.96 | 32.40 | 31.48 | 34.16 | 42.49 | 27.89 | 35.99 | 31.20 | 26.51 | 24.10 | 32.00 | 28.45 | 24.19 | 29.20 | 30.80 | 3.92 | 26.55 | 32.48 |
| Dispider [13] | 1 FPS | 74.92 | 75.53 | 74.10 | 73.08 | 74.44 | 59.92 | 76.14 | 62.91 | 62.16 | 45.80 | 67.63 | 35.46 | 25.26 | 38.57 | 43.34 | 35.66 | 39.62 | 27.65 | 34.80 | 25.34 | 33.61 | 53.12 |
| Models under StreamBridge (Offline Streaming), Multi-Turn Evaluation | |||||||||||||||||||||||
| Oryx-1.5-7B† [1] | 1 FPS | 78.47 | 77.17 | 83.86 | 80.20 | 71.07 | 66.98 | 79.63 | 61.38 | 66.29 | 40.93 | 70.59 | 30.00 | 15.20 | 33.60 | 43.20 | 30.50 | 20.40 | 24.80 | 39.60 | 54.90 | 34.93 | 53.76 |
| + Stream-IT | 1 FPS | 82.29 | 77.95 | 87.98 | 86.47 | 77.99 | 81.31 | 76.85 | 69.92 | 71.96 | 35.23 | 74.79 | 19.20 | 14.40 | 52.00 | 29.20 | 28.70 | 14.40 | 14.80 | 51.20 | 43.14 | 30.89 | 54.79 |
| LLaVA-OV-7B† [3] | 1 FPS | 76.84 | 77.17 | 82.60 | 75.25 | 64.15 | 64.17 | 75.00 | 61.38 | 61.19 | 46.11 | 68.39 | 24.40 | 12.00 | 32.40 | 37.60 | 26.60 | 20.00 | 19.60 | 34.40 | 52.94 | 31.74 | 50.96 |
| + Stream-IT | 1 FPS | 82.29 | 72.44 | 92.09 | 80.86 | 71.07 | 74.46 | 75.00 | 62.20 | 70.26 | 28.50 | 70.92 | 20.80 | 13.20 | 43.60 | 27.60 | 26.30 | 20.40 | 17.60 | 41.60 | 37.26 | 29.21 | 51.73 |
| Qwen2-VL-7B† [2] | 1 FPS | 80.38 | 78.74 | 83.22 | 79.86 | 74.21 | 69.47 | 77.78 | 63.41 | 69.97 | 43.01 | 72.01 | 32.00 | 15.20 | 39.60 | 38.40 | 31.30 | 25.20 | 21.20 | 33.20 | 66.67 | 36.57 | 55.09 |
| + Stream-IT | 1 FPS | 84.74 | 82.68 | 88.92 | 89.77 | 77.36 | 85.36 | 84.26 | 69.92 | 71.67 | 35.75 | 77.04 | 18.00 | 13.20 | 43.60 | 21.60 | 24.10 | 14.00 | 17.20 | 48.00 | 50.98 | 32.55 | 55.39 |
Appendix F Limitations
While our proposed framework and dataset significantly enhance the streaming capabilities of existing offline Video-LLMs, there are still limitations worth noting. First, although the activation model is compact and decoupled from the main LLM, its decision quality is inherently limited by the feature granularity of the small activation model’s visual encoder. In highly complex or ambiguous scenes, the model may still struggle to accurately predict the optimal response timing. Second, while Stream-IT provides large-scale multi-turn, interleaved training data, its construction relies partially on synthetic QA generation and clip concatenation, which, despite careful filtering, may introduce domain shift compared to truly continuous, real-world video streams. Future work could benefit from curating more organically collected long-form streaming videos with naturally evolving events and dialogues. Third, although our round-decayed compression improves inference efficiency and memory usage, it is based on frame-level pruning. Further investigation is needed to develop more adaptive compression strategies that dynamically balance historical preservation and token budget based on scene complexity at a token level. Finally, StreamBridge currently focuses on frame-by-frame streaming under relatively low sampling rates (e.g., 1 FPS). Extending the framework to handle denser frame rates or multi-modal streaming inputs (e.g., audio-visual-text) in real-time remains an important direction for future research.

### v2

@@
 In contrast to prior efforts [10; 11; 12; 13], which train streaming models from scratch but fall behind on offline video tasks, StreamBridge leverages the strong generalization capabilities of existing Video-LLMs without requiring full retraining.
-Concretely, StreamBridge introduces a memory buffer to manage incoming video frames, coupled with a round-decayed compression strategy that merges earlier frame tokens while preserving recent ones, enabling the model to support long-context, multi-turn interactions in streaming scenarios.
+This approach allows developers to directly benefit from the rich world knowledge and linguistic fluency of large-scale pre-trained models, while incurring only minimal additional computational cost and data requirements.
+Concretely, StreamBridge introduces a memory buffer to manage incoming video frames, coupled with a round-decayed compression strategy that merges earlier frame tokens while preserving recent ones, enabling the model to support long-context, multi-modal, and multi-turn interactions in streaming scenarios.
 For proactive capabilities, instead of modifying the base model architecture [14] or introducing streaming-specific objectives [12], both of which can lead to optimization conflicts and issues like probability correction [10], StreamBridge adopts a modular design, by decoupling the proactive capability from the main Video-LLM via a compact activation model.
@@
 Our method introduces a compact, plug-and-play activation model that enables proactive behaviors efficiently and non-intrusively.
+We also provide additional discussions on how our work relates to the ReKV [56], VideoStreaming [57], and StreamChat [53] in the Appendix F.
 3 Methodology
@@
 This process is encapsulated in the compression function in Algorithm 1 (line 15).
-The detailed pseudo codes can be found in Appendix G.
+The detailed pseudo codes can be found in Appendix I.
 3.2.3 A Plug-and-play Activation Model
-To enable proactive responses, we decouple the activation function from the main Video-LLM.
+To enable proactive responses in streaming Video-LLMs, we decouple the activation function from the main Video-LLM.
 Unlike prior methods that tightly integrate activation mechanisms into the LLM [10; 12; 14; 52], our framework avoids potential interference with the language modeling capacity of the main Video-LLM.
@@
 This design enables the model to learn temporal dependencies and identify appropriate response moments throughout the video stream.
-For training data, we collect a diverse set of temporally annotated video datasets across multiple tasks, including dense video captioning [57; 58], sequential step recognition [59; 60], grounded video question answering [61; 62; 63], temporal video grounding [64], and temporal action detection [65; 66].
+For training data, we collect a diverse set of temporally annotated video datasets across multiple tasks, including dense video captioning [59; 60], sequential step recognition [61; 62], grounded video question answering [63; 64; 65], temporal video grounding [66], and temporal action detection [67; 68].
 For each task, we design specific prompt templates and randomly select among them as during training (details in Appendix A).
@@
 As analyzed in Section 3.1, streaming scenarios are primarily characterized by multi-turn real-time understanding and proactive responses.
-However, existing datasets and video sources fall short of fully supporting these requirements [67; 68].
+However, existing datasets and video sources fall short of fully supporting these requirements [69; 70].
 To fill this gap and further enhance the streaming interaction capability of StreamBridge, we introduce Stream-IT—a video-text dataset specifically designed for streaming instruction tuning with an interleaved multi-turn dialogue format.
 Datasets for Proactive Understanding.
-We collect a set of public datasets enriched with timestamp annotations, spanning a wide range of tasks including: (i) Dense Video Captioning [57; 58; 69]; (ii) Sequential Step Recognition [59; 60; 70]; (iii) Grounded VideoQA [61; 71; 72; 73].
+We collect a set of public datasets enriched with timestamp annotations, spanning a wide range of tasks including: (i) Dense Video Captioning [59; 60; 71]; (ii) Sequential Step Recognition [61; 62; 72]; (iii) Grounded VideoQA [63; 73; 74; 75].
 All datasets are reformatted into a proactive-style interleaved format: ‘<> <> <>, <> <>, ’, where may be an open-ended query (e.g., “Who is the man going to find?”) or a goal-oriented instruction (e.g., “Show me all the steps for cooking.”).
-Unlike traditional single-turn datasets where a question is immediately followed by an answer [67; 68], our structure introduces a temporal delay between <> and <> through the inserted video segments <>, simulating proactive response scenarios.
+Unlike traditional single-turn datasets where a question is immediately followed by an answer [69; 70], our structure introduces a temporal delay between <> and <> through the inserted video segments <>, simulating proactive response scenarios.
 StreamingQA-120K: Multi-Turn, Long-Form QA Construction.
@@
 General Video Understanding.
-While our StreamBridge is designed for online scenarios, we also verify that it does not downgrade the base model’s performance on standard offline video tasks.
+While our method is designed for online scenarios, we also verify that it does not downgrade the base model’s performance on standard offline video tasks.
 As shown in Table 2, models equipped with the StreamBridge framework and fine-tuned on Stream-IT (denoted with ‡) exhibit consistent improvements or maintain comparable performance relative to their original versions.
@@
 Finally, removing the synthetic StreamingQA-120K subset from Stream-IT degrades performance across both streaming and offline benchmarks, underscoring the crucial role of StreamingQA-120K in boosting both streaming and offline video understanding capabilities.
+| OVO-Bench | VideoMME | |
+| (Real-Time) Avg. | Avg. | |
+| 4k | 70.49 | 61.7 |
+| 8k | 70.89 | 63.6 |
+| 16k | 71.30 | 64.4 |
+| 32k | 71.16 | 64.7 |
+Impact of .
+To better understand the impact of , we conducted ablation studies using the Qwen2-VL-StreamBridge model with 1 FPS sampling, varying from 4k to 32k.
+From Table 6, we observe the following: (1) For streaming tasks (e.g., OVO-Bench Real-Time): Model performance remains relatively stable across varying values, ranging from 70.49% to 71.30%; Accuracy peaks at 16k and slightly declines at 32k, suggesting that further increasing the memory budget yields diminishing returns.
+This supports our design assumption: in streaming scenarios, models primarily rely on recent context, and older frames can be compressed without significant performance loss.
+(2) For offline tasks (e.g., VideoMME): Accuracy improves consistently as increases, from 61.7% at 4k to 64.7% at 32k.
+This means that offline tasks benefit more from retaining the full temporal context and more uncompressed video tokens, especially for long videos that require detailed long-range understanding.
+StreamBridge can flexibly balance efficiency and performance across both streaming and offline settings by adjusting the memory budget accordingly, and we set = 16k to strike a good balance between them across most tasks.
 Activation Threshold.
@@
 The procedure is repeated using as the new anchor, generating from , and so on.
-This results in a similarity-ordered list of videos .
+This results in a similarity-ordered list of videos We formulate this process in Algorithm 2.
 This approach allows for flexible concatenation of any number of clips to construct a long-form sample, by directly selecting a continuous span , without re-computing similarity each time.
@@
 For evaluation metrics, we compute the average F1 score across multiple IoU thresholds (IoU {}) for localization-based tasks.
-For tasks involving text generation, we adopt sentence-level similarity metrics [83] to measure the semantic alignment between model outputs and ground-truth responses, following prior works [64; 13].
+For tasks involving text generation, we adopt sentence-level similarity metrics [85] to measure the semantic alignment between model outputs and ground-truth responses, following prior works [66; 13].
 Specifically, the all-MiniLM-L6-v2 model in Sentence-Transformers library is used as the embedding model.
@@
 Moreover, the results of TVGF1, TALF1, are the same for our method with different main Video-LLMs, since they use the same activation model and will not be affected by the generated response.
-Appendix E Full Results on OVO-Bench and Streaming-Bench
+Appendix E Broader Impacts
+There are many real-world applications of streaming Video-LLMs, such as patient or elderly health monitoring, autonomous driving, and collaborative robots.
+However, there could be unintended usages and we advocate responsible usage complying with applicable laws and regulations.
+Appendix F More Related Works
+To address the challenge of long-context understanding in streaming video, several memory and retrieval mechanisms have been proposed.
+For instance, ReKV [56] introduces a training-free framework that stores and retrieves the Key-Value (KV) caches of processed frames, enabling offline models to answer user queries efficiently by reloading only the most relevant context.
+Besides, VideoStreaming [57] employs a memory-propagated encoding architecture where a condensed representation of the preceding clip serves as historical context for encoding the next, combined with an adaptive selection of memories for question-answering.
+Moreover, StreamChat [53] proposes a hierarchical memory system comprising short-term, long-term, and dialogue components to facilitate complex streaming interactions, and also contributes the StreamBench benchmark for evaluating diverse streaming scenarios.
+While these methods effectively advance long-context retention for reactive question-answering, StreamBridge differs by introducing a round-decayed compression strategy specifically tailored for multi-turn real-time interactions, which efficiently prunes redundant historical tokens while preserving recent context with high fidelity.
+Moreover, StreamBridge introduces a decoupled, lightweight activation model.
+This plug-and-play component operates in parallel with the main Video-LLM, enabling continuous proactive responses.
+These designs, supported by our dedicated Stream-IT dataset, effectively transform general-purpose offline models into versatile and proactive streaming assistants without compromising their core performance.
+Appendix G Limitations
+Although our proposed framework and dataset significantly enhance the streaming capabilities of existing offline Video-LLMs, there are still limitations worth noting.
+First, while Stream-IT provides large-scale multi-turn, interleaved training data, its construction relies partially on synthetic QA generation and clip concatenation, which, despite careful filtering, may introduce domain shift compared to truly continuous, real-world video streams.
+Future work could benefit from curating more organically collected long-form streaming videos with naturally evolving events and dialogues.
+Second, StreamBridge currently focuses on frame-by-frame streaming under relatively low sampling rates (e.g., 1 FPS).
+Extending the framework to handle denser frame rates or multi-modal streaming inputs (e.g., audio-visual-text) in real-time remains an important direction for future research.
+Appendix H Full Results on OVO-Bench and Streaming-Bench
 | Method | # of | Real-Time Visual Perception | Backward Tracing | Forward Active Responding | Overall. | ||||||||||||
@@
 | + Stream-IT | 1 FPS | 84.74 | 82.68 | 88.92 | 89.77 | 77.36 | 85.36 | 84.26 | 69.92 | 71.67 | 35.75 | 77.04 | 18.00 | 13.20 | 43.60 | 21.60 | 24.10 | 14.00 | 17.20 | 48.00 | 50.98 | 32.55 | 55.39 |
-Appendix F Limitations
-While our proposed framework and dataset significantly enhance the streaming capabilities of existing offline Video-LLMs, there are still limitations worth noting.
-First, although the activation model is compact and decoupled from the main LLM, its decision quality is inherently limited by the feature granularity of the small activation model’s visual encoder.
-In highly complex or ambiguous scenes, the model may still struggle to accurately predict the optimal response timing.
-Second, while Stream-IT provides large-scale multi-turn, interleaved training data, its construction relies partially on synthetic QA generation and clip concatenation, which, despite careful filtering, may introduce domain shift compared to truly continuous, real-world video streams.
-Future work could benefit from curating more organically collected long-form streaming videos with naturally evolving events and dialogues.
-Third, although our round-decayed compression improves inference efficiency and memory usage, it is based on frame-level pruning.
-Further investigation is needed to develop more adaptive compression strategies that dynamically balance historical preservation and token budget based on scene complexity at a token level.
-Finally, StreamBridge currently focuses on frame-by-frame streaming under relatively low sampling rates (e.g., 1 FPS).
-Extending the framework to handle denser frame rates or multi-modal streaming inputs (e.g., audio-visual-text) in real-time remains an important direction for future research.